# IBIS: A Powerful Hybrid Architecture for Human Activity Recognition

### v1

@@ III-D Support Vector Machine with Radial Basis Function Kernel
IV Performance Evaluation
@@
This section evaluates the hybrid IBIS neural network using generalization data from Scenario S7. This scenario, a laboratory room filled with desks, chairs, and monitors, was chosen as a benchmark due to its challenging environment. The numerous obstacles in this setup introduce complex signal distortions like multipath propagation, reflection, and scattering, which make the task of activity recognition particularly difficult for any classifier. We will analyze the model’s performance through its learning curves, class-wise confusion matrices, and the computational cost of the hybrid architecture.
IV-A Confusion Matrix for 5 Movements
@@
Figure 4 compares the confusion matrices for five movements, showing the results from the Inception-only network against the proposed hybrid scheme. As seen in Figure 4a, the Inception network struggled with precision, particularly for the walking movement, which was frequently confused with running, achieving only 71% accuracy. Jumping was also often misclassified as running, with an accuracy of just 67%. This lack of precision is attributed to the similarity of Doppler patterns between these movements and the fact that the Inception network was not designed to handle temporal dependencies crucial for accurate CSI data classification.
@@
Figure 4: Comparative analysis of the confusion matrix for five movements: original Inception versus the hybrid IBIS with SVM post-processing.
@@
Conversely, Figure 4b shows the results after implementing our hybrid IBIS model and the SVM algorithm, which led to significantly better performance with only minor misclassifications. The improvement is due to the BiLSTM component’s ability to extract temporal patterns from the Doppler data. Furthermore, the SVM, with its non-linear kernel, proper C parameter tuning, and class balancing, refined the model’s classifications, leading to enhanced overall performance. Both confusion matrices presented refer to the post-antenna fusion phase, where classification decisions are combined via majority voting.
IV-B ROC Curve for 5 Movements
@@
Figure 5: ROC curve analysis for five movements: comparing the original Inception with the hybrid Neural Network IBIS with SVM Post-Processing.
@@
As shown in Figure 5a, the original model performed poorly on the walking, running, and jumping classes, which showed significant fluctuations in accuracy. In contrast, Figure 5b demonstrates that with our proposed approach, accuracy values improved and stabilized with no apparent oscillations, even without antenna fusion. This highlights the effectiveness of applying ensemble learning in this challenging scenario.
IV-C Confusion Matrix for 8 Movements
@@
Figure 6 shows the confusion matrices for eight movements, including: standing, stand up, and arm gym. In Figure 6a, the original algorithm’s performance worsened compared to the five-movement scenario, with a higher misclassification rate for gestures like arm gym, jumping, and sitting. This is likely due to the SHARP algorithm’s sanitization process, where the probability of error increases proportionally with the number of classified gestures and bandwidth [12]. This behavior indicates that the original model struggled with overfitting and failed to generalize when more movements were added.
@@
Figure 6: Comparative analysis of the confusion matrix for eight movements: original Inception versus the hybrid IBIS with SVM post-processing.
@@
In contrast, Figure 6b demonstrates that the overall accuracy remained high despite the addition of new movements. This superior performance is attributed to the combination of the IBIS network and the SVM algorithm. Even after attempting to improve the original model’s performance by modifying the dataset split to 80% for training, 10% for testing, and 10% for validation, the results remained unchanged.
IV-D ROC Curve for 8 Movements
Figure 7 presents the ROC curves for all classes in the eight-movement scenario. The original simulation, shown in Figure 7a, performed poorly on the running and jumping classes, which exhibited constant oscillations and lower accuracy. This behavior is attributed to the model’s inability to generalize effectively, resulting in misclassifications and a higher incidence of false positives.
Figure 7: ROC curve analysis for eight movements: comparing the original Inception with the hybrid IBIS approach with SVM Post-Processing.
On the other hand, Figure 7b demonstrates significantly improved results with our proposed method. The curves show no apparent oscillations and maximum accuracy was achieved across all classes. These results confirm that combining the Inception-BiLSTM architecture with the SVM algorithm is highly effective, as the performance was considerably enhanced even without antenna fusion.
IV-E Comparative Algorithms
@@
Table III demonstrates that the hybrid Inception-BiLSTM + SVM model achieved the best and most consistent results compared to other ensemble learning models. Its superior performance is attributed to the neural network’s ability to capture complex patterns and the SVM’s capacity to handle temporal sequence boundaries, both of which are essential for Doppler-based data. The original Inception model, for example, struggled with classifying similar movements like walking and jumping, leading to false positives and poor generalization.
TABLE III: Models Accuracy
@@
IBIS+SVM 100 100 98.69 97.41 97.53
Inception 100 100 70.79 100 66.67
Note: SVM - Support Vector Machine
Similarly, Table IV, which presents average metrics for five movement classes, confirms that our IBIS + SVM model outperformed other tested models. We included additional metrics like F1-score, recall, and precision to provide a more comprehensive evaluation, as relying solely on accuracy can be misleading. These metrics help assess the model’s performance in imbalanced scenarios and its ability to balance precision and recall.
TABLE IV: Average Models Metrics
@@
IBIS+SVM 98.95 98.78 98.72 98.85
Inception 88.12 88.01 87.49 91.52
Note: SVM - Support Vector Machine
Our method also showed a significant improvement over a transfer learning approach using a hybrid CNN-ABLSTM model from a related study[8]. Our proposed IBIS + SVM achieved an accuracy of 98.95% compared to their 85.82%, a remarkable 18.62% improvement under the same condition of 100 training epochs. This is because our approach uses a combination of signal processing (the SHARP algorithm for noise elimination), deep learning (the enhanced hybrid network for feature and temporal learning), and traditional machine learning (the SVM for refining classification). This pipeline and the inclusion of more attention layers, increased architectural complexity, and Doppler traces features, significantly enhanced generalization compared to the other study, which used raw unprocessed data.
IV-F Computational Complexity
@@
  • An Inception network combined with a Bi-directional Long Short-Term Memory (BiLSTM) layer.
The standalone Inception network had the lowest computational cost and the shortest training time per epoch due to its shallow architecture. However, it failed to accurately classify movements, producing false positives and showing temporal instability.
V Conclusion
This study successfully addresses the challenge of Human Activity Recognition (HAR) using Wi-Fi sensing, a non-intrusive and low-cost technology with broad applications in healthcare, surveillance, and monitoring. We introduced a novel hybrid neural network architecture, IBIS, that combines Inception-BiLSTM with a SVM to enhance data generalization and classification accuracy. Our model’s performance was rigorously evaluated across scenarios with five and eight activities, achieving an exceptional average accuracy of nearly 99%. This not only demonstrates the method’s robustness against noise and interference but also represents a significant improvement of approximately 10.94% over the state-of-the-art SHARP study. The excellent results were further validated by high-performing ROC curves and other metrics.
We also conducted a comparative analysis, confirming that the Inception-BiLSTM + SVM combination outperformed other configurations. This underscores the effectiveness of our approach, which strategically leverages convolutional layers for feature extraction, BiLSTM for temporal analysis, and SVM for refining decision boundaries. For future work, we propose exploring refinements in the Doppler curves acquisition process, so that the complexity of the neural network training/post-processing can be reduced.

### v2

@@ 3.2 Post-Processing Support Vector Machine
3.3 Pseudocode
The pseudocode description presented in algorithm  1 concisely illustrates the application of the IBIS framework and the main processes involved.
It is worth mentioning that the code containing the modifications to the SHARP algorithm, as well as all stages of the IBIS implementation, will be made public after the publication of this proposal.
Algorithm 1 IBIS framework
1:Preprocessing:
2:    Apply LASSO-based sanitization and remove offset
3:    Extract Doppler traces and select subcarriers
4:Input: Dataset $X$, labels $Y$, batch size 3232, learning rate $0.001$, epochs 100100
5:    Normalize features: $X_{\text{norm}}=\text{normalize}(X)$
6:    Split dataset: $(X_{train},Y_{train},X_{test},Y_{test})\leftarrow\text{split}(X_{\text{norm}},Y,0.8)$
7:Inception Feature Extraction:
8:    Apply parallel convolutions with multiple kernel sizes
9:    Concatenate outputs to obtain $F_{\text{inception}}$
10:BiLSTM Temporal Modeling:
11:    Feed $F_{\text{inception}}$ into BiLSTM layers
12:    Extract temporal representation $F_{\text{BiLSTM}}$
13:Softmax Classification:
14:    Compute posterior probabilities:
15:     $P(Y|X)=\text{Softmax}(W_{s}F_{\text{BiLSTM}}+b_{s})$
16:    Perform Grid Search to select kernel type, $C$, and $\gamma$
17:SVM Training (Postprocessing):
18:    Train SVM using $(P(Y|X_{train}),Y_{train})$
19:    Available kernels:
20:     RBF: $K(x,x^{\prime})=\exp(-\gamma\|x-x^{\prime}\|^{2})$
21:     Sigmoid: $K(x,x^{\prime})=\tanh(\gamma\,x^{\top}x^{\prime}+r)$
22:     Polynomial: $K(x,x^{\prime})=(\gamma\,x^{\top}x^{\prime}+r)^{d}$
23:Prediction:
24:    Compute $P(Y|X_{test})$ using the Inception–BiLSTM–Softmax network
25:    Predict: $Y_{pred}=\text{SVM.predict}(P(Y|X_{test}))$
26:Evaluation:
27:    Compute Accuracy, Precision, Recall, and F1-score
28:Output: Final predictions $Y_{pred}$ and evaluation metrics
4 Performance Evaluation
@@
This section evaluates the IBIS framework using generalization data from Scenario S7 for Dataset A and from Scenario S2 for Dataset B, as previously described. It is important to note that this validation used an 80 MHz bandwidth (IEEE 802.11ac) for Scenario S7 and a 160 MHz bandwidth (IEEE 802.11ax) for Scenario S2, with the selection of four antennas to maximize the amount of information obtained during training. This scenario, a laboratory room filled with desks, chairs, and monitors, was chosen as a benchmark due to its challenging nature. The numerous obstacles in this setup introduce complex signal distortions like multipath propagation, reflection, and scattering, which make the task of activity recognition particularly difficult for any classifier. We will analyze the model’s performance through its learning curves, class-wise confusion matrices, and the computational cost of the hybrid architecture.
For data separation, after Doppler extraction, the dataset was divided into 60%, 20%, and 20% for training, testing, and validation, respectively.
To avoid decision bias in the evaluated algorithms, ten simulations were performed for each experiment, and the results provided herein represent the average across the ten simulations.
4.1 Confusion Matrix for 5 Movements
@@
Figure 4 compares the confusion matrices for five movements, showing the results from the Inception-only network against the proposed hybrid scheme. As seen in Figure 4a, the Inception network struggled with precision, particularly for the walking movement, which was frequently confused with running, achieving only 71% accuracy. Jumping was also often misclassified as running, with an accuracy of just 70%. This lack of precision is attributed to the similarity of Doppler patterns between these movements and the fact that the Inception network was not designed to handle temporal dependencies crucial for accurate CSI data classification.
(a) Inception - 88.16%
(b) IBIS - 95.40%
Note: E - Empty, S - Sitting, W - Walking, R - Running, and J - Jumping
@@
Figure 4: Comparative analysis of the confusion matrix for five movements: original Inception versus IBIS
Conversely, Figure 4b shows the results obtained after implementing our IBIS framework, which achieved substantially better performance with only minor misclassifications. A more balanced distribution of accuracy across the classes is also observed, reaching up to 99% in categories where the original model previously exhibited confusion. Overall, this configuration yields a 7.58% performance gain compared to the previous simulation.
@@
This improvement is primarily attributed to the BiLSTM component’s ability to capture temporal patterns from the Doppler data. In addition, the SVM—through the selection of an appropriate kernel, proper tuning of the parameter $C$, and class balancing—further refined the classification boundaries, resulting in enhanced overall performance. Both confusion matrices shown correspond to the post-antenna fusion stage, where classification decisions are merged using majority voting.
4.2 ROC Curve for 5 Movements
@@
Figure 5: ROC curve analysis for five movements: comparing the original Inception with IBIS
@@
As shown in Figure 5a, the original model performed poorly for the walking, running, and jumping classes, which exhibited significant fluctuations in accuracy. In contrast, Figure 5b shows that, with our proposed approach, the accuracy values increased and became more stable, with no apparent oscillations, even without antenna fusion, except for the Walking class. At this stage, we are evaluating only the effect of applying the SVM after training the neural network using the softmax probabilities as input. In other words, this result reflects exclusively the effectiveness of the SVM refinement, excluding the majority-voting step from antenna fusion, which is applied afterward. However, when considering the class distribution, the results are more uniform. This highlights the effectiveness of incorporating ensemble learning in this challenging scenario.
4.3 Confusion Matrix for 8 Movements
@@
Figure 6 shows the confusion matrices for eight movements, including standing, stand up, and arm gym. In Figure 6a, the performance of the original algorithm deteriorated compared to the five-movement scenario, exhibiting a higher misclassification rate for gestures such as arm gym, jumping, and sitting, resulting in an accuracy of 38.38%. This is likely due to the SHARP algorithm’s sanitization process, where the error probability increases proportionally to the number of classified gestures and bandwidth [9]. This behavior indicates that the original model struggled with overfitting and failed to generalize when more movements were added.
(a) Inception - 38.38%
(b) IBIS - 70.25%
Note: E - Empty, S - Sitting, W - Walking, R - Running, J - Jumping, U - Stand Up, G - Standing and A - Arm Gym
@@
Figure 6: Comparative analysis of the confusion matrix for eight movements: original Inception versus IBIS
@@
In contrast, Figure 6b shows that the overall accuracy remained high despite the addition of new movements, reaching 70.25%. This superior performance is attributed to the combination of the Inception-BiLSTM hybrid network and the SVM algorithm. Even after attempting to improve the original model’s performance by modifying the dataset split to 80% for training, 10% for testing, and 10% for validation, the results remained unchanged.
4.4 Doppler Traces
In Figure 7, we present the Doppler trace plots for eight activities. It is possible to observe that the traces corresponding to the Empty, Sitting, Standing, and Arm Gym activities exhibit very similar movement patterns, making them difficult for a neural network to distinguish. Consequently, their spatial features lie very close to one another. In this context, the BiLSTM network combined with the SVM becomes the decisive component: the BiLSTM is able to extract temporal dependencies, and once these dependencies are captured, the SVM refines the decision boundaries through appropriate tuning of the trade-off parameter, kernel choice, and margin.
It becomes clear that, even with IBIS, the framework struggles to reliably distinguish highly similar Doppler traces due to the strong overlap between activities. This reinforces the need for a carefully designed preprocessing pipeline. Moreover, depending on the normalization strategy, relevant offsets and peak-related features may be suppressed, which are essential for accurate neural network inference. Nevertheless, IBIS still outperforms the original Inception network, yielding an improvement of approximately 45.30%.
Figure 7: Doppler traces for the activities: E – Empty, S – Sitting, W – Walking, R – Running, J – Jumping, U – Stand Up, G – Standing, A – Arm Gym.
4.5 Ablation
The next experiment conducted with our framework aims to evaluate its behavior when individual components of the pipeline are removed. First, we run the experiment using the SVM together with the Inception network and the full IBIS model. In the second step, we remove the SVM from the composition, leaving only the Inception–BiLSTM neural network. Analyzing the first experiment in Figure 8, we observe that the Inception model with SVM experiences a reduction in performance, achieving 85.27%, whereas the full IBIS model reaches 95.40%.
Removing the SVM from the pipeline resulted in a 2% increase in accuracy, highlighting the contribution of the BiLSTM component, which achieved 90% compared to the 88% produced by the original Inception network. When this performance is compared with the 95.40% obtained by the complete framework, it becomes clear that the SVM plays a predominant role, as most of the accuracy gain originates from its inclusion in the post-processing stage, where it effectively refines class boundaries with high precision.
Another experiment involved removing the Doppler traces from the pipeline and applying IBIS directly to the sanitized data. Analyzing Figure 8, on the right-hand side, we observe that the Inception network without Doppler information reaches 34% accuracy, while IBIS achieves 49%. Although both results are lower, this confirms that Doppler features play a fundamental role in the analysis of phase components. For this specific dataset, removing Doppler information leads to a substantial degradation in accuracy, especially for the Inception network.
Figure 8: Graph illustrating the behavior of the accuracy when removing the Doppler traces and/or the SVM components
4.6 Exploring another dataset
With the intention of demonstrating the efficiency of the IBIS framework, new experiments were conducted using the public dataset provided by Cominelli et al. [9]. In this experiment, Scenario S2 was used, and—similarly to the previous experiments—the Grid Search technique was applied to select the trade-off parameter $C$ and the $\gamma$ parameter, as well as to evaluate different kernels, including RBF, Sigmoid, and Polynomial. Thus, the exact same procedure adopted in our framework was applied here, but now with a different dataset.
In this previously described dataset, several SHARP adaptations were applied to support operation with a 160 MHz bandwidth. Five activities—Walking, Running, Jumping, Sitting, and Empty Room, were evaluated, as shown in the confusion matrix presented in Figure 9. Based on this dataset, adjustments were also made to the scripts to make them more consistent with those used in the other dataset.
It is important to highlight that, unlike the previous dataset, the IBIS framework was not able to effectively distinguish between the walking and running activities in this dataset. This limitation is mainly due to the reduced amount of available information: after generating the Doppler traces, the window segments created during the dataset splitting and training stages become significantly smaller. As a result, the IBIS does not receive sufficient information to reliably differentiate between these two activities.
Analyzing Figure 9, we observe that the Inception network alone achieved an accuracy of 70.23%, whereas the IBIS framework reached 81.40%, representing an accuracy gain of 13.72%. This difference in performance between datasets can be attributed to variations in data distribution; elements such as obstacles, walls, and interference introduce additional complexity and reduce recognition performance due to the increased similarity between activity patterns.
These experiments demonstrate that IBIS maintains superior performance across different datasets, achieving substantial improvements even under varying environmental conditions and inherent limitations.
(a) Inception - 70.23%
(b) IBIS - 81.40%
Note: A - Walking, B - Run, C - Jump, D - Sitting, and E - Empty Room
Figure 9: Comparative analysis of the confusion matrices for five activities: original Inception model versus the IBIS, using the dataset from Cominelli et al. [9]
4.7 Comparable Algorithms
@@
Table 3 demonstrates that the IBIS framework achieved the best and most consistent results compared to other ensemble learning models. Its superior performance is attributed to the neural network’s ability to capture complex patterns and the SVM’s capacity to handle temporal sequence boundaries, both of which are essential for Doppler-based data. The original Inception model, for example, struggled with classifying similar movements like walking and jumping, leading to false positives and poor generalization.
Table 3: Models Accuracy
@@
IBIS 100 87 91 100 99
Inception 100 99 71 100 70
Similarly, Table 4, which presents average metrics for five movement classes, confirms that our IBIS outperformed other tested models. We included additional metrics like F1-score, recall, and precision to provide a more comprehensive evaluation, as relying solely on accuracy can be misleading. These metrics help assess the model’s performance in imbalanced scenarios and its ability to balance precision and recall.
Table 4: Average Models Metrics for 5 Movements
@@
IBIS 95.40 95.21 95.30 95.57
Inception 88.16 88.26 87.87 91.47
While the CNN-ABLSTM model from a related study [10] presents a lower computational cost due to its lighter architecture, it achieves only 85.82% accuracy, whereas our IBIS method reaches 95.40%. This 10.04% improvement under the same 100-epoch training condition is primarily due to the limited number of attention layers and the reduced architectural complexity in their model, which hinder its ability to capture complex temporal–spatial patterns and subtle Doppler-induced variations. In contrast, our IBIS framework leverages the Inception module for spatial feature extraction and the BiLSTM module for temporal dependency learning, making it more robust to noise, interference, and reflective-surface effects. Additionally, our pipeline integrates advanced signal processing, the SHARP algorithm for noise suppression, and traditional machine learning SVM for refined classification. This combined approach significantly improves generalization compared to the transfer-learning-based model trained on raw, unprocessed data in  [10].
4.8 Computational Complexity
@@
  • An Inception network combined with a BiLSTM layer.
The standalone Inception network had the lowest computational cost and the shortest training time per epoch due to its shallow architecture. However, it failed to accurately classify movements, producing false positives and showing temporal instability. It should be noted that the SVM classifier is not displayed in this figure because it is not involved in the epoch-level training process. The SVM is fitted only once after the neural network has converged, using the extracted feature probabilities. Consequently, its computational overhead is negligible, and its inclusion in the graphical representation is unnecessary.
5 Conclusion
This paper addresses the challenges of generalization and robustness in HAR using non-intrusive Wi-Fi sensing. We propose IBIS, a novel hybrid architecture that strategically integrates Inception layers for spatial feature extraction, a BiLSTM module for temporal modeling, and an SVM classifier for sharper decision boundaries. Extensive evaluations across scenarios with different activity sets (five and eight classes) demonstrate the framework’s effectiveness, reaching an average accuracy of 95.40%. Notably, IBIS outperforms the state-of-the-art SHARP method by 7.58% on Dataset A, highlighting its superior capability in handling noise and environmental interference. A second improvement of 13.40% was observed on Dataset B, largely due to environmental differences, where variations in surface materials and layout can alter signal propagation patterns. Nevertheless, even under new conditions and datasets, IBIS consistently delivers superior performance. The analysis further confirms that replacing the standard Softmax layer with an SVM classifier significantly enhances classification performance in complex environments. Future work will focus on optimizing the Doppler trace acquisition and pre-processing stages to reduce computational complexity and further streamline the training pipeline.